Q. India plays two ODI matches each with Australia and Pakistan. The probability of India getting 0, 1, 2 are 0.45, 0.05, 0.50. The probability of India getting at least 7 points in the series is

A) 0.00875
B) 0.875
C) 0.0875
D) None of these
Explanation

(c)
Matches played by India = 4
Maximum points in any match = 2
$\therefore$ Maximum points in four matches can be 8 only.
Therefore, atleast 7 points means 7 or 8 points.
$\therefore\$ Required probability $= P(7) + P(8)$
$= {\ ^{4}C}_{1}(0.05)(0.5)^{3} + (0.5)^{4}$
$= 0.0250 + 0.0625 = 0.0875$

Dr. Abubacker
Reviewed by Dr. Abubacker

Physics Educator & NEET/JEE Specialist

Ph.D. in Physics with 19+ years of teaching experience.